OpenStudy (anonymous):

Two factory plants are making TV panels. Yesterday, Plant A produced 4 times as many panels as Plant B. Five percent of the panels from Plant A and 4% of the panels from Plant B were defective. How many panels did Plant A produce, if the two plants together produced 1440 defective panels?

OpenStudy (anonymous):

let call x= amount of panels produced by PlantB it's means that Plant A produces 4x(4 times) right?

OpenStudy (anonymous):

yes right

OpenStudy (anonymous):

5% from Plant A = 0.05*4x 4% from Plant B = 0.04*x this is amounts of defective panels, right?

OpenStudy (anonymous):

i would assume so. yes

OpenStudy (anonymous):

We know that it's equal 1440 let's put equation together: 0.05*4x+0.04x=1440 now - solve for x

OpenStudy (anonymous):

k hang on! i can do it:)

OpenStudy (anonymous):

i'll wait

OpenStudy (anonymous):

or u can go ahead. i wont look lol

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

0.2x + 0.04x = 1440 0.24x=1440 x=1440/0.24 x=144000/24 x=6000

OpenStudy (anonymous):

ok sweet!!! thank you very very much! i was going brain dead on where to even start! thank you

OpenStudy (anonymous):

remember, that x is x= amount of panels produced by PlantB to find production of Plant A - multiply by 4 (4 times as many panels as Plant B) 6000*4=..... :)) Good Luck & Welcome!